OpenStudy (anonymous):

Use graphing to find the solutions to the system of equations. y=x^2+7x+7 y=x+2

OpenStudy (mathmath333):

so u are allowed to use graphing calculator

OpenStudy (mathmath333):

https://www.desmos.com/calculator/1rk2cnk3m5

OpenStudy (mathmath333):

otherwise \(\large \begin{align} \color{black}{ y=x^2+7x+7\\ y=x+2 \hspace{.33em}\\~\\ x^2+7x+7=x+2 \hspace{.33em}\\~\\ x^2+6x+5=0 \hspace{.33em}\\~\\ x^2+5x+x+5=0 \hspace{.33em}\\~\\ x^2+5x+x+5=0 \hspace{.33em}\\~\\ x(x+5)+1(x+5)=0 \hspace{.33em}\\~\\ (x+1)(x+5)=0 \hspace{.33em}\\~\\ x=-1~~or~~x=-5 \hspace{.33em}\\~\\ }\end{align}\) put each value of x to get y in original equations